# Self-Enhancing Video Data Management System for Compositional Events with Large Language Models

### v1

@@ Abstract.
Complex video queries can be answered by decomposing them into modular subtasks. However, existing video data management systems assume the existence of predefined modules for each subtask. We introduce VOCAL-UDF, a novel self-enhancing system that supports compositional queries over videos without the need for predefined modules. VOCAL-UDF automatically identifies and constructs missing modules and encapsulates them as user-defined functions (UDFs), thus expanding its querying capabilities. To achieve this, we formulate a unified UDF model that leverages large language models (LLMs) to aid in new UDF generation. VOCAL-UDF handles a wide range of concepts by supporting both program-based UDFs (i.e., Python functions generated by LLMs) and distilled-model UDFs (lightweight vision models distilled from strong pretrained models). To resolve the inherent ambiguity in user intent, VOCAL-UDF generates multiple candidate UDFs and uses active learning to efficiently select the best one. With the self-enhancing capability, VOCAL-UDF significantly improves query performance across three video datasets.
@@ 1. Introduction
The rapid advancement of video analytics has fueled the development of innovative applications across various fields. In medical education, surgery videos enhance students’ procedural knowledge by illustrating complex temporal and spatial events (Friedl et al., 2006). In biology, scientists use wildlife footage to study organism behaviors and interactions in their natural habitats (SCH, 2013; Dellinger et al., 2021). In transportation, traffic surveillance systems analyze and manage traffic flow, improving urban mobility (Hammer et al., 2020). Across these applications, analysts seek to query video databases for events characterized by spatio-temporal and semantic interactions. For instance, an analyst might search for “a motorcycle swerving near a silver Subaru and then colliding with it” or “a doctor holding a scalpel and then placing it on a table.”
@@
Though promising, answering video queries using frontier vision-language models remains underwhelming. Although vision-language models have demonstrated notable capabilities on diverse, challenging tasks (Radford et al., 2021; Liu et al., 2023; Li et al., 2023b), they struggle to answer compositional queries (Ma et al., 2023) that involve combinations of recognizing objects (e.g., “car”, “truck”), reasoning about relationships (e.g., “behind”, “holding”), and identifying attributes (e.g., “silver color”, “Subaru make”). This challenge is further amplified when queries require temporal reasoning (Yu et al., 2023) (e.g., “X then Y”, “X for at least 10 seconds”). While new models continue to improve their ability to reason spatially (Chen et al., 2024) and compositionaly (Hu et al., 2023), their performance remains low (Wang et al., 2024). Additionally, deploying large models at scale is prohibitively expensive (Qiao et al., 2024) and inference is slow (Wan et al., 2023). For example, current vision-language models can only achieve a throughput of around 100 tokens per second (Bai et al., 2023; gpt, 2024), making their use in large-scale video analytics intractable.
@@
Alternatively, workflow-oriented video data management systems (VDBMSs) answer compositional queries by decomposing them into granular subtasks (Andreas et al., 2016; Gupta and Kembhavi, 2023; Surís et al., 2023; Shen et al., 2023; Fu et al., 2019; Bastani et al., 2020b; Zhang et al., 2023; Mell et al., 2023). Submodules identify objects in frames, their locations, attributes, relationships, and track them across frames. These individual elements—objects, attributes, and relationships—form spatio-temporal scene graphs (Krishna et al., 2017; Ji et al., 2020), which can expressively representation many complex visual queries. Various scene graph generation techniques (Zellers et al., 2018; Dai et al., 2017) have been proposed to automatically extract scene graphs from images and videos. In those systems, the subtasks, which extract scene graph elements, are first solved individually (e.g., a system might leverage a relationship classifier to identify relationships between objects in a frame) and are then composed to answer a compositional query (Lu et al., 2016).
VDBMSs rest on a critical assumption: the existence of modules capable of executing the subtasks that comprise a complex query. Systems typically predefined a variety of modules to support its use cases (Gupta and Kembhavi, 2023; Surís et al., 2023; Shen et al., 2023; Zhang et al., 2023; Mell et al., 2023) and many support extensibility via user-defined functions (UDFs) to address unsupported scenarios (Zhang et al., 2023; Xu et al., 2022; Kakkar et al., 2024). For example, to answer our initial motorcycle query, a system might support built-in identification for cars but a user might need to supply a UDF to filter for “silver” objects or determine when objects are “near”.
@@
Despite the availability of pre-trained computer vision models that can be readily integrated as UDFs, there are many instances where users require solutions for domain-specific applications or seek to identify fine-grained object classes and subjective concepts for which no off-the-shelf models exist. Identifying or adapting domain-specific models in such cases may be possible, though tedious. Additionally, it may be necessary to—on a per-UDF basis—curate datasets and perform extensive training to achieve satisfactory performance (Andreas et al., 2016).
@@
To address these challenges, we present VOCAL-UDF (Figure 1), a new self-enhancing video data management system that empowers users to flexibly issue and answer compositional queries, even when the modules necessary to answer those queries are unavailable. To use VOCAL-UDF, a user only needs to provide a video dataset and a natural language (NL) description of the query. VOCAL-UDF then automatically identifies and builds the necessary modules and encapsulates them as new UDFs to expand its querying capabilities. It then compiles the NL query into a declarative one that it executes over the video dataset.
@@
There are several challenges in building VOCAL-UDF. First, given an NL query, VOCAL-UDF converts it into a declarative one that it can execute. Unfortunately, existing methods assume the existence of predefined modules that can be invoked to construct the declarative query (Zhang et al., 2023; Daum et al., 2022). The important challenge that VOCAL-UDF addresses is the identification, during the compilation process, of any semantic concepts that are not supported by existing modules or UDFs. In such scenarios, new UDFs need to be created. To address this challenge, VOCAL-UDF leverages an LLM’s reasoning capabilities to determine when (and which) new UDFs are needed.
@@
The second challenge involves transforming the identified missing concepts into executable modules. Video queries depend on a wide variety of semantic concepts and VOCAL-UDF must generate UDFs that can accurately classify them (e.g., ”near” or ”holding”). While LLMs have been shown to produce useful code in various contexts (Chen et al., 2021; Austin et al., 2021; Ye et al., 2023; Chen et al., 2023a), queries on video data are often highly ambiguous and their performance varies significantly (see Section 6.3). To address this challenge, VOCAL-UDF automatically implements two types of UDFs—program-based UDFs and distilled-model UDFs—enabling it to handle a wide range of concepts. Program-based UDFs, which take the form of imperative Python functions in our system, operate on relational tables and video pixels. This class of UDFs can already classify many relationships and attributes with high quality (Zhang et al., 2023; Chen et al., 2019). However, we show in Section 4.3 the need for distilled-model UDFs, which are lightweight ML models that are trained on the fly to classify more nuanced concepts (Toubal et al., 2024). For both classes, VOCAL-UDF again leverages LLMs: when generating program-based UDFs, VOCAL-UDF uses an LLM’s programming capability, whereas distilled-model UDF implementations rely on the LLM’s ability to annotate and convert visual concepts into machine learning (ML) models (Toubal et al., 2024). Additionally, because an LLM may generate incorrect or erroneous UDFs in both cases, VOCAL-UDF implements syntax and semantic verification steps.
@@
The third challenge involves managing the inherit ambiguity in user intent when articulating a query. This is especially important for specialized and subjective concepts that are difficult to resolve without user feedback. For example, users may have different interpretations of the “near” relationship. Additionally, in different contexts either a program-based or distilled-model UDF maybe produce the best results. Finally, for the most challenging semantic concepts, no UDF may perform well, risking the overall query performance. To address this combination of challenges, rather than producing a single, one-size-fits-all UDF to match a semantic concept, VOCAL-UDF instead generates multiple candidate UDFs with different semantic interpretations, implementations, and properties. VOCAL-UDF then employs active learning to efficiently identify the implementation that best matches the user’s intent.
@@
Finally, the self-building nature of our system necessitates a consistent and unified UDF model. Typically, a UDF can be an arbitrary function that operates on database tuples. However, one challenge lies in modeling UDFs in a structured way to handle concepts of objects, relationships, and attributes while seamlessly integrating and interacting with various system components. To address this, we formulate a unified UDF scheme for different semantic concepts, which enables the LLM to generate UDFs in a structured format and simplifies the compilation process. By incrementally growing the database, future UDFs can be constructed from existing ones.
In summary, VOCAL-UDF makes the following contributions:
We propose a unified scheme of UDFs for objects, relationships, and attributes in videos (Section 3).
We propose a self-building framework that leverages LLMs to automatically expand VOCAL-UDF’s querying capabilities over time. To do so, VOCAL-UDF parses NL queries and turns unseen semantic concepts into either program-based UDFs or distilled-model UDFs (Sections 4.2 and 4.3).
We develop a method to effectively manage both the ambiguity in semantic concepts and the error-prone nature of LLMs to generate a diverse set of candidate UDFs and efficiently determine the one that best aligns with the user’s intent (Section 4.4).
@@
We evaluate VOCAL-UDF on three video datasets from different domains (Yi et al., 2020; Feng et al., 2021; Ji et al., 2020) and show that it significantly improves query performance, in terms of F1 score, by automatically selecting, implementing, and executing its automatically-generated UDFs (Section 6).
@@
Overall, thanks to its self-enhancing capability, VOCAL-UDF is an important step toward making VDBMSs more practical to deploy and use in a variety of applications.
@@ 2. Background
VOCAL-UDF adopts the scene graphs data model and query language presented in EQUI-VOCAL (Zhang et al., 2023), which models the compositional video events as spatio-temporal scene graphs. This approach draws from cognitive foundations in human perception (Zacks et al., 2001; Kurby and Zacks, 2008; Biederman, 1987) and has been shown to enable a variety of compositional queries (Zhang et al., 2023). This section summarizes key background information about these concepts.
@@
DSL. VOCAL-UDF adopts EQUI-VOCAL’s domain-specific language (DSL) (Zhang et al., 2023). This DSL encapsulates the logical structure of a query while abstracting away the full details inherent to SQL. EQUI-VOCAL’s query executor compiles these DSL queries into SQL, allowing them to be executed over relational tables. As an example, the query “A car is initially far from a truck, then moves close to the truck for more than 10 seconds” might be expressed as (assuming a frame rate of 24 frames per second):
In this DSL, the variable represents an arbitrary object in a query, with distinct subscripts indicating objects with different ’s. All predicates in a region graph are separated by commas. Region graphs are then sequenced in temporal order using semicolons. Each region graph can persist for multiple frames and there can be other frames between two adjacent region graphs. Finally, the notation is employed to stipulate that the region graph exists in at least consecutive frames. The query returns a set of video segment identifiers.
@@ 3. A UDF-Based Data Model
In this section, we formalize the types of UDFs supported by VOCAL-UDF (Section 3.1). These UDFs allow users to define custom objects, relationships, and attributes. We then describe how UDFs are compiled and executed in VOCAL-UDF (Section 3.2).
3.1. An extensible video-UDF framework
@@
Table 1 shows the relational schema of VOCAL-UDF. The Frames relation includes a virtual column pixels that stores the pixel values of each video frame in a 3D array with dimensions , where , , and are the height, width, and the three color channels of the frame. The Objects, Relationships, and Attributes relations store the detected objects, relationships, and attributes in the videos, respectively.
A user-defined function (UDF) is a function that the user creates and a database engine executes to extend the functionality of the system. A typical UDF input is zero or more columns, and it returns a scalar value or a row set per invocation. Once created, a UDF can be used in expressions in SQL statements, e.g., in WHERE clauses.
@@
When querying video databases, VOCAL-UDF supports executing and generating UDFs that can be used to identify custom objects, relationships, and attributes in videos. Including UDFs in the query statement enables users to find complex, compositional events in videos. LABEL:lst:sql_snippet shows an example SQL query with UDFs for the objects car and truck, the relationship behind, and the attribute silver. VOCAL-UDF supports imperative and declarative UDFs; our implementation uses Python. UDFs in VOCAL-UDF are categorized into the following classes:
A relationship UDF or attribute UDF is a scalar, predicate function that classifies whether a pair of objects (for relationships) or an object (for attributes) exhibits the specified relationship or attribute. It operates on domains of existing columns and/or frame pixels to identify a new relationship or attribute. The UDF accepts zero (i.e., a dummy UDF; see Section 4.4.2) or more columns as input and returns a boolean value. The input arguments can include any of the following columns from each table in the FROM clause of the SQL query: pixels, oname, x1, y1, x2, y2, rname, and aname.
[fontsize=]python bool rel_or_attr_udf(a1, a2, …, an)
VOCAL-UDF restricts UDFs to be frame-level, i.e., they operate on object(s) within the same frame. Therefore, input arguments (a1, a2, …, an) are all from the same video frame (identified by vid and fid) and are associated with one or two distinct objects.
As an example, we can define a relationship UDF to check whether object is behind another object by comparing their centroid -coordinates. {minted}[fontsize=]python def behind(o1_y1, o1_y2, o2_y1, o2_y2): return (o1_y1 + o1_y2) / 2 ¡ (o2_y1 + o2_y2) / 2 We can also define an attribute UDF to check whether a detected car is silver by running an ML model over the crop of the video frame pixels: {minted}[fontsize=]python def silver(pixels, x1, y1, x2, y2): cropped_img = pixels[y1:y2, x1:x2] is_silver = awesome_color_classifier(cropped_img) return is_silver
@@
An Object UDF requires localizing, classifying, and tracking objects in videos. Instead of returning a boolean value, it is a table-valued function that takes a video segment video_pixels, as input. A video segment comprises frames with the same vid concatenated into a 4D array with an additional dimension for the frame index. An object UDF detects and tracks objects of a specific class. Given a video segment, the UDF makes calls to a custom object detection and tracking model and returns a row set, which follows the Objects schema listing the detected and tracked objects in the video segment. {minted}[escapeinside=——, fontsize=]python —table— object_udf(video_pixels)
As a concrete example, we can define an object UDF that detects all cars in a video: {minted}[fontsize=]python def car(video_pixels): obj_tuples = [] car_detector, tracker = load_models() for frame in video_pixels: detected_cars = car_detector(frame) tracked_objs = tracker.update(detected_cars) obj_tuples.extend(tracked_objs) return obj_tuples
@@
A Value-lookup UDF is a class of UDFs that simply encapsulate a predicate over existing column values. {minted}[fontsize=]python bool value_udf(a1) As an example, suppose the value “car” is in the domain of the object oname. Then, a value-lookup UDF can be defined as: {minted}[fontsize=]python def car(oname): return oname == ’car’
@@
3.2. Compilation and optimization.
@@
Object, relationship, and attribute UDFs can be expensive to evaluate, since they can operate on image pixels and invoke ML models. To optimize query execution, VOCAL-UDF caches the results of invoking expensive UDFs and replaces them with value-lookup UDFs. When executing any object, relationship, or attribute UDF over the video corpus for the first time, VOCAL-UDF automatically materializes the results to make it available as a value in the corresponding column and replaces the UDF with a value-lookup UDF with the same function name. If an object has multiple attribute entries or a pair of objects has multiple relationship entries, multiple tuples will be generated and added to the table. For example, when running a query with the silver attribute UDF for the first time, VOCAL-UDF evaluates the UDF for each object in the video corpus. If an object OID1 in a frame F1 of a video segment V1 is classified as silver, VOCAL-UDF inserts a new row with values (V1, F1, OID1, ‘silver’) into the Attributes relation. Then, VOCAL-UDF replaces the silver attribute UDF with a value-lookup UDF of the same function name that checks whether the aname is ‘silver’. In our implementation, we materialize the results by executing the UDF over the entire dataset. Executing UDFs that invoke ML models can be accelerated through batch inference. Reusing the results of predicate evaluation for query optimization is a long-standing research topic (Xu et al., 2022; Roy et al., 2000; Mistry et al., 2001; Elghandour and Aboulnaga, 2012), which is not the focus of this paper.
4. Proposing and generating new UDFs
To use VOCAL-UDF, users provide a video dataset and optionally a set of UDFs. Users can then issue natural language queries to identify compositional events within these videos. VOCAL-UDF translates these textual queries into our DSL notation using the available UDFs and then executes queries to find matching video segments. When needed, VOCAL-UDF automatically generates new UDFs to process the query.
Following the same example, the traffic engineer loads the traffic surveillance videos into VOCAL-UDF and provides an object detection and tracking model that can detect and track cars and trucks as the UDF. To effectively answer the user’s query, VOCAL-UDF can automatically propose and generate new UDFs to classify “near”, “silver color”, and “Subaru make” and use them together with the existing UDF to execute the query.
Figure 2 shows the architecture of VOCAL-UDF. The Query Parser takes a user query described in natural language and a set of available UDFs as input and parses the query into the DSL notation described in Section 2. If successful, the DSL query is passed to the Query Executor for execution over the video dataset to find all matching videos. If the query contains predicates that cannot be resolved using available UDFs, the UDF Generator is invoked to build new UDFs. The Query Parser can then re-parse the query with the updated UDFs. The Storage Manager maintains the raw videos, extracted relational information, and available UDFs.

### v2

@@ Abstract.
Complex video queries can be answered by decomposing them into modular subtasks. However, existing video data management systems assume the existence of predefined modules for each subtask. We introduce VOCAL-UDF 111The source code is available at https://github.com/uwdb/VOCAL-UDF, a novel self-enhancing system that supports compositional queries over videos without the need for predefined modules. VOCAL-UDF automatically identifies and constructs missing modules and encapsulates them as user-defined functions (UDFs), thus expanding its querying capabilities. To achieve this, we formulate a unified UDF model that leverages large language models (LLMs) to aid in new UDF generation. VOCAL-UDF handles a wide range of concepts by supporting both program-based UDFs (i.e., Python functions generated by LLMs) and distilled-model UDFs (lightweight vision models distilled from strong pretrained models). To resolve the inherent ambiguity in user intent, VOCAL-UDF generates multiple candidate UDFs and uses active learning to efficiently select the best one. With the self-enhancing capability, VOCAL-UDF significantly improves query performance across three video datasets.
@@ 1. Introduction
Rapid advances in video analytics have fueled the development of innovative applications across various fields. In medical education, surgery videos enhance students’ procedural knowledge by illustrating complex temporal and spatial events (Friedl et al., 2006). In biology, scientists use wildlife footage to study organism behaviors and interactions in their natural habitats (Schlining et al., 2013; Dellinger et al., 2021). In transportation, traffic surveillance systems analyze and manage traffic flow, improving urban mobility (Hammer et al., 2020). Across these applications, analysts seek to query video databases for events characterized by spatio-temporal and semantic interactions. For instance, an analyst might search for “a motorcycle swerving near a silver Subaru and then colliding with it” or “a doctor holding a scalpel and then placing it on a table.”
@@
Though promising, answering video queries using frontier vision-language models (VLMs) remains underwhelming. Although VLMs have demonstrated notable capabilities on diverse, challenging tasks (Radford et al., 2021; Liu et al., 2023; Li et al., 2023b), they struggle to answer compositional queries (Ma et al., 2023) that involve recognizing objects (e.g., “car”, “truck”), reasoning about relationships (e.g., “behind”, “holding”), and identifying attributes (e.g., “silver color”, “Subaru make”). This challenge is further amplified when queries require temporal reasoning (Yu et al., 2023) (e.g., “X then Y”, “X for at least 10 seconds”). While new models continue to improve their ability to reason spatially (Chen et al., 2024d) and compositionally (Hu et al., 2023), their performance remains low (Wang et al., 2024). Additionally, deploying large models at scale is prohibitively expensive (Qiao et al., 2024) and inference is slow (Wan et al., 2024). For example, current VLMs can only achieve a throughput of around 100 tokens per second (Bai et al., 2023; gpt, 2024), making their use in large-scale video analytics intractable.
@@
Alternatively, workflow-oriented video data management systems (VDBMSs) answer compositional queries by decomposing them into granular subtasks (Andreas et al., 2016; Gupta and Kembhavi, 2023; Surís et al., 2023; Shen et al., 2023; Fu et al., 2019; Bastani et al., 2020b; Zhang et al., 2023b; Mell et al., 2023). Submodules identify and track objects, attributes, and relationships across frames, forming spatio-temporal scene graphs (Krishna et al., 2017; Ji et al., 2020). These scene graphs can expressively represent many complex visual queries. Various scene graph generation techniques (Zellers et al., 2018; Dai et al., 2017) have been proposed to extract scene graphs from images and videos. In these systems, subtasks that extract scene graph elements are solved individually and then composed to answer a compositional query (Lu et al., 2016).
@@
However, VDBMSs make a critical assumption: the existence of modules capable of executing subtasks to answer a complex query. Systems typically provide a variety of built-in modules (Gupta and Kembhavi, 2023; Surís et al., 2023; Shen et al., 2023; Zhang et al., 2023b; Mell et al., 2023) and often allow extensibility via user-defined functions (UDFs) for unsupported scenarios (Zhang et al., 2023b; Xu et al., 2022; Kakkar et al., 2024). Concerning our motorcycle query, a user might need to supply a UDF to filter for “silver” or nearby objects if the system lacks these capabilities. Despite the availability of pre-trained computer vision models that can be readily integrated as UDFs, users may require solutions for domain-specific applications or seek to identify fine-grained object classes and subjective concepts for which no off-the-shelf models exist. Identifying or adapting domain-specific models in such cases may be possible, though tedious. Additionally, it may be necessary to—on a per-UDF basis—curate datasets and perform extensive training to achieve satisfactory performance (Andreas et al., 2016).
@@
To address these challenges, we present VOCAL-UDF (Figure 1), a self-enhancing VDBMS that empowers users to flexibly issue and answer compositional queries, even when the necessary modules are unavailable. To use VOCAL-UDF, a user only needs to provide a video dataset and a natural language (NL) description of the query. VOCAL-UDF then automatically identifies and builds the necessary modules and encapsulates them as new UDFs to expand its querying capabilities. It then compiles the NL query into a declarative one that it executes over the video dataset.
@@
There are several challenges in building VOCAL-UDF. First, given an NL query, VOCAL-UDF converts it into a declarative one that it can execute. Unfortunately, existing methods assume the existence of predefined modules that can be invoked to construct the declarative query (Zhang et al., 2023b; Daum et al., 2022). VOCAL-UDF addresses this challenge by identifying any semantic concepts that are not supported by existing modules or UDFs during compilation and leveraging an LLM’s reasoning to determine when and which new UDFs are needed.
@@
The second challenge involves transforming the various identified missing semantic concepts into executable modules. While LLMs can produce useful code in various contexts (Chen et al., 2021; Austin et al., 2021; Ye et al., 2023; Chen et al., 2024a), queries on video data are often highly ambiguous and their performance varies significantly (see Section 6.3). To address the challenge of handling a multitude of missing concepts, VOCAL-UDF automatically implements two types of UDFs—program-based UDFs and distilled-model UDFs. In our system, program-based UDFs are imperative Python functions that operate on relational tables and video pixels. While this class of UDFs can classify many relationships and attributes with high quality (Zhang et al., 2023b; Krishna et al., 2019), we show in Section 4.3 the need for distilled-model UDFs, which are lightweight machine learning (ML) models that are trained on the fly to classify more nuanced concepts (Toubal et al., 2024). VOCAL-UDF leverages LLMs for both UDF types: program-based UDFs harness an LLM’s programming capability, whereas distilled-model UDFs rely on the LLM’s ability to annotate and convert visual concepts into ML models (Toubal et al., 2024). To mitigate potential erroneous LLM-generated UDFs, VOCAL-UDF implements syntax and semantic verification steps.
@@
The third challenge involves managing the inherent ambiguity in user intent when articulating a query. This is especially important for specialized and subjective concepts that are difficult to resolve without user feedback. For example, users may have different interpretations of the “near” relationship. Further compounding this difficulty is our observation that for some inputs a program-based UDF might perform better than a distilled-model UDF, while in other cases the opposite is true. Finally, for the most challenging semantic concepts, no UDF may perform well and employing one might risk overall query performance. To address this formidable combination of challenges, VOCAL-UDF generates multiple candidate UDFs each with different semantic interpretations, implementations, and properties. VOCAL-UDF then employs active learning to efficiently identify the variant that best matches the user’s intent.
@@
Finally, the self-building nature of our system necessitates a consistent and unified UDF model. Typically, a UDF can be an arbitrary function that operates on database tuples. However, one challenge lies in modeling UDFs in a structured way to handle concepts of objects, relationships, and attributes while seamlessly integrating and interacting with various system components. To address this, we propose a unified UDF scheme for different semantic concepts, which enables the LLM to generate UDFs in a structured format and simplifies the compilation process. By incrementally growing the database, future UDFs can be composed using existing ones.
@@
We evaluate VOCAL-UDF on three video datasets from different domains (Yi et al., 2020; Feng et al., 2021; Ji et al., 2020) and show that it significantly improves query performance, in terms of F1 score, by automatically selecting, implementing, and executing its automatically-generated UDFs. We finally conduct a thorough analysis of VOCAL-UDF to understand its efficiency in terms of execution time and monetary cost.
@@
Overall, VOCAL-UDF’s self-enhancing capability is an important step toward making VDBMSs more practical to deploy and use in a variety of applications.
@@ 2. Background
VOCAL-UDF leverages EQUI-VOCAL’s (Zhang et al., 2023b) scene graph data model and query language, which models compositional video events as spatio-temporal scene graphs. This approach draws from cognitive foundations in human perception (Zacks et al., 2001; Kurby and Zacks, 2008; Biederman, 1987) and enables a variety of compositional queries (Zhang et al., 2023b). This section summarizes key background information about these concepts.
@@
Query language. In VOCAL-UDF, a query identifies video segments that match a user-specified event. VOCAL-UDF supports relational queries over the schema in Table 1, which are of the following form. Using Datalog notation: , where is a temporally ordered sequence of region graphs specifying that a matching event consists of , followed by , followed by , etc. Each can persist for multiple frames and there can be other frames between and . p is a set of predicates that are applied to objects, relationships, and attributes in region graphs. d is a set of constraints on the duration for which a region graph must remain valid before transitioning to the next one. Lastly, is the maximum number of frames between and . As an example, the event “A car is initially far from a truck, then remains close to the truck for more than 10 seconds” can be expressed as in LABEL:lst:datalog_snippet.
DSL. VOCAL-UDF adopts EQUI-VOCAL’s domain-specific language (DSL) (Zhang et al., 2023b), which encapsulates query logic while abstracting away SQL details. The query executor compiles these DSL queries into efficient SQL for execution over relational tables. Using the DSL, the same event can be expressed as (assuming 24 frames per second): . In this DSL, the variable represents an arbitrary object in a query, with distinct subscripts indicating objects with different ’s. All predicates in a region graph are separated by commas. Region graphs are then sequenced in temporal order using semicolons. Each region graph can persist for multiple frames and there can be other frames between two adjacent region graphs. Finally, stipulates that the region graph exists in at least consecutive frames. The query returns a set of video segment identifiers.
@@ 3. A UDF-based data model
In this section, we formalize the types of UDFs supported by VOCAL-UDF. These UDFs allow users to define custom objects, relationships, and attributes. We then describe how UDFs are compiled and executed in VOCAL-UDF.
@@
Table 1 shows VOCAL-UDF’s relational schema. The Frames relation includes a virtual pixels column that stores the frame pixel values in a 3D array (, where is height, is width, and 3 represents the color channels). The Objects, Relationships, and Attributes relations store corresponding detected frame elements.
@@
A UDF extends database functionality. A typical UDF takes columns as input, returns a scalar value or a row set, and is used in SQL statements, e.g., in WHERE clauses. When querying video databases, VOCAL-UDF supports generating and executing UDFs to identify custom objects, relationships, and attributes, enabling users to find complex, compositional events. LABEL:lst:sql_snippet shows an example SQL query with UDFs for the objects car and truck, the relationship behind, and the attribute silver. VOCAL-UDF supports imperative and declarative Python-based UDFs, and are categorized into the following classes: relationship, attribute, object, and value-lookup.
A relationship UDF or attribute UDF is a deterministic, scalar predicate that indicates whether an input exhibits a specified relationship or attribute. It accepts zero (i.e., a dummy UDF; see Section 4.4.2) or more columns as arguments and produces a boolean result. The parameters can include any of the following columns from each table in the SQL query’s FROM clause: pixels, oname, x1, y1, x2, y2, rname, and aname. VOCAL-UDF restricts relationship UDFs and attribute UDFs to be frame-level, i.e., they operate on object(s) within the same frame. Therefore, input arguments are all from the same video frame (identified by vid and fid) and are associated with one or two distinct objects. As an example, we might define a relationship UDF to indicate whether object is behind another object by comparing their centroid -coordinates: {minted}[fontsize=]python def behind(o1_y1, o1_y2, o2_y1, o2_y2): return (o1_y1 + o1_y2) / 2 ¡ (o2_y1 + o2_y2) / 2 We could also define an attribute UDF to indicate if a detected car is silver by running an ML model over the frame pixels: {minted}[fontsize=]python def silver(pixels, x1, y1, x2, y2): cropped_img = pixels[y1:y2, x1:x2] is_silver = awesome_color_classifier(cropped_img) return is_silver
@@
An object UDF requires localizing, classifying, and tracking objects in videos. Instead of returning a boolean value, it is a table-valued function that takes a video segment, video_pixels, as input. A video segment comprises frames with the same vid concatenated into a 4D array with an additional dimension for the frame index. An object UDF detects and tracks objects of a specific class. Given a video segment, the UDF makes calls to a custom object detection and tracking model and returns a row set, which follows the Objects schema listing the detected and tracked objects. As an example, we can define an object UDF that detects all cars in a video: {minted}[fontsize=]python def car(video_pixels): obj_tuples = [] car_detector, tracker = load_models() for frame in video_pixels: detected_cars = car_detector(frame) tracked_objs = tracker.update(detected_cars) obj_tuples.extend(tracked_objs) return obj_tuples
@@
A value-lookup UDF is a class of UDFs that simply encapsulate a predicate over existing column values. As an example, suppose the value “car” is in the domain of the object oname. Then, a value-lookup UDF can be defined as: {minted}[fontsize=]python def car(oname): return oname == ’car’
@@
| ObjView(vid, videopixels) |
| RelView(vid, fid, pixels, rid, o1o2rnames, o2o1rnames, |
| o1oid, o1x1, o1y1, o1x2, o1y2, o1anames, |
| o2oid, o2x1, o2x2, o2y1, o2y2, o2anames) |
| AttrView(vid, fid, pixels, oid, oname, x1, y1, x2, y2, anames) |
UDFs can have different lists of parameters in their signatures. To execute a query with UDFs, VOCAL-UDF first constructs the relational views as shown in Table 2 derived from the relations in Table 1. The views–ObjView, RelView and AttrView–contain all attributes that an object UDF, relationship UDF, and attribute UDF can potentially accept as input arguments, respectively. For simplicity, VOCAL-UDF generates UDFs with a uniform list of parameters for each class.
@@
Object, relationship, and attribute UDFs can be expensive to evaluate, as they may operate on image pixels and invoke ML models. To optimize query execution, VOCAL-UDF caches the results of all UDFs and replaces them with value-lookup UDFs. When executing a UDF over a video corpus for the first time, VOCAL-UDF materializes the results to make it available as a value in the corresponding column and substitutes the UDF with a value-lookup UDF of the same name. For example, when running a query with the silver attribute UDF for the first time, VOCAL-UDF evaluates the UDF for each object in the video corpus. If an object OID1 in a frame F1 of a video segment V1 is classified as silver, VOCAL-UDF inserts a new row with values (V1, F1, OID1, ‘silver’) into the Attributes relation. Then, VOCAL-UDF replaces the silver attribute UDF with a value-lookup UDF that checks whether the aname is ‘silver’. Reusing the results of predicate evaluation for query optimization is a long-standing research topic (Xu et al., 2022; Roy et al., 2000; Mistry et al., 2001; Elghandour and Aboulnaga, 2012) and is not the focus of this paper. Batch inference can further accelerate ML-based UDFs.
4. VOCAL-UDF approach